OpenStudy (anonymous):

On a science test, 2 points are deducted from a total of 100 points for each question answered incorrectly. Trevor scored an 84 on the test. The equation 100 – 2x = 84, where x represents the number of questions answered incorrectly, represents the situation. How many questions did Trevor answer incorrectly

OpenStudy (isaidavila):

Simply solve the equation to find x that is the number of incorrect answer 100-2x=84 -2x=-16 2x=16 x=8